# Pure Vision Language Action (VLA) Models: A Comprehensive Survey

### v2

@@ Abstract
The emergence of Vision Language Action (VLA) models marks a paradigm shift from traditional policy-based control to generalized robotics, reframing Vision Language Models (VLMs) from passive sequence generators into active agents for manipulation and decision-making in complex, dynamic environments. This survey delves into advanced VLA methods, aiming to provide a clear taxonomy and a systematic, comprehensive review of existing research. It presents a comprehensive analysis of VLA applications across different scenarios and classifies VLA approaches into several paradigms: autoregression-based, diffusion-based, reinforcement-based, hybrid, and specialized methods; while examining their motivations, core strategies, and implementations in detail. In addition, foundational datasets, benchmarks, and simulation platforms are introduced. Building on the current VLA landscape, the review further proposes perspectives on key challenges and future directions to advance research in VLA models and generalizable robotics. By synthesizing insights from over three hundred recent studies, this survey maps the contours of this rapidly evolving field and highlights the opportunities and challenges that will shape the development of scalable, general-purpose VLA methods.
@@
Vision Language Action, Vision Language Model, Robotics, Embodied AI
@@ 1 Introductions
To address these limitations, researchers have begun exploring the incorporation of large language models (LLMs) and vision language models (VLMs) to enable more accurate and flexible robotic manipulation. Modern robotic manipulation methods [1, 2] typically leverage vision language generative paradigms (e.g., autoregressive models [3, 4, 5, 6] or diffusion models [7]), combined with large-scale datasets [8] and advanced fine-tuning strategies. We refer to these as VLA foundation models, which have substantially improved the quality of robotic manipulations. Fine-grained action control over generated content provides users with greater flexibility, unlocking the practical potential of VLA for task execution.
@@
Figure 3: Vision Language Action Taxonomy: From Autoregression-based, Diffusion-based, to Reinforcement-based and Hybrid/Specialized methods, multi-paradigm advances and practical adaptations in VLA. The taxonomy is organized along a chronological timeline.
@@ 2 Backgrounds
The emergence of Vision Language Action (VLA) models represents a significant advance toward general-purpose embodied intelligence. Traditional robotic systems typically rely on isolated perception pipelines, hand-engineered control strategies, or task-specific reinforcement learning. Although these approaches perform effectively in constrained environments, such as factory floors or laboratories, they generalize poorly to dynamic and unstructured settings. Modern robots can “see” using computer vision models, “understand” language through large language models, and “act” via controllers or learned policies; however, integrating these capabilities into a coherent and unified system remains a key challenge. VLA models address this challenge by providing a unified framework that grounds language in perception and maps it to executable actions.
@@ 2.1 Early: LLM/VLM fundamental Models
Against this backdrop, vision language models (VLMs) emerged as a crucial bridge between unimodal learning and embodied intelligence. Early approaches (e.g., ViLBERT [16], VisualBERT [17]) aligned and fused images and text using dual-stream or single-stream Transformers, while contrastive learning methods (e.g., CLIP) mapped large-scale image–text pairs into a shared embedding space, enabling zero-shot and few-shot recognition and retrieval. More recently, instruction-tuned, dialogue-centric multimodal models (e.g., BLIP-2 [18], Flamingo [19], LLaVA [20]) have substantially strengthened open-ended cross-modal understanding, fine-grained grounding, and multi-turn reasoning, setting the stage for vision language action (VLA) systems.
@@ 2.2.1 From LLM/VLM to VLA Models
Following this trajectory, research has naturally advanced toward VLA integration, which unifies visual perception, language understanding, and executable control within a single sequence modeling framework[21, 22, 1, 2]. Typical designs encode images and instructions as prefix or context tokens, inject robot states and sensory feedback as state tokens, and autoregressively generate action tokens to produce control sequences, thereby closing the perception-language-action loop. Compared with traditional perception, planning, and control pipelines, VLA provides end-to-end cross-modal alignment and a unified treatment of goals, constraints, and intent. It inherits the semantic and instruction generalization of VLMs, while explicit state coupling and action generation confer robustness to environmental disturbances and long-horizon tasks. This progression—from unimodal to multimodal, and then to multimodal plus executable control—establishes the methodological foundation for systems that can not only see and understand, but also act.
@@ 2.2.2 The Supporting Role of Data and Simulation
The development of vision language action (VLA) models in robotics relies heavily on high-quality datasets and realistic simulators capable of capturing the complexity of real-world scenarios. Recent robotic methods are typically deep learning–based and data-driven; consequently, dataset collection and annotation play a critical role in driving advancements in the field. Some datasets are collected in real-world settings, which requires substantial human effort and financial resources. To address these challenges, researchers also leverage large-scale human manipulation videos from the Internet as generalization datasets to provide auxiliary supervision for VLA model training. Despite these efforts, data collection remains costly, annotations are labor-intensive, and long-tail corner cases are often underrepresented. Other datasets are generated through robot simulators, which facilitate the collection of large-scale labeled data. Simulators provide diverse and controllable environments, flexible sensor configurations, realistic kinematic models, and interactive static and dynamic scenes, supporting both data collection and model evaluation. Representative datasets include Open X-Embodiment (OXE) [8], which integrates 22 robot datasets from 21 institutions, encompassing 527 skills and 160,266 tasks, and BridgeData [23], comprising 71 tasks across 10 environments in multiple domains. These resources standardize data formats, thereby promoting rapid development and reproducibility in VLA research. Simulators such as THOR [24], Habitat [25], MuJoCo [26], Isaac Gym [27], and CARLA [28] offer extensible virtual environments capable of generating multimodal annotations, including action trajectories, object states, and natural language instructions. Collectively, these datasets and simulation platforms mitigate the scarcity of real-world robot data and accelerate the training and evaluation of VLA models.
@@ 2.3 Future: Towards General Embodied Intelligence
VLA models occupy the forefront of research where vision, language, and action converge. They build upon breakthroughs in foundational models for perception and reasoning, emphasizing capabilities in human interaction and task execution, and extend these abilities to the physical world. By integrating the representational power of visual encoders, the reasoning capabilities of large language models, and the decision-making abilities of reinforcement learning and control frameworks, VLA models hold significant potential to bridge the “perception–understanding–action” gap. Despite facing challenges related to scalability, generalization, safety, and real-world deployment, VLA is widely recognized as a key frontier in embodied artificial intelligence. Although VLA has achieved notable success in vision language action interactions, and benefits from advances in large-scale language models, it has not yet attained full generality in the domain of embodied intelligence. General embodied intelligence posits that human-like intelligent behavior relies not only on cognitive processing but also on the physical body, environmental perception, and feedback mechanisms, enabling interaction with the external world. To accommodate the demands of diverse tasks, general embodied intelligence can manifest through various types of robots, including humanoid robots for household applications, assembly robots with dexterous manipulators, and bionic robots with specialized capabilities. Clearly, general embodied intelligence has the potential to enable artificial systems to perform a wider array of tasks across diverse environments. VLA is currently evolving toward this vision of general embodied intelligence and holds considerable promise for realizing it.
3 Vision Language Action Models
@@
In recent years, Vision Language Action (VLA) models have undergone rapid and systematic development, driven by advances in multimodal representation learning, generative modeling, and reinforcement learning. To trace this evolution, this section reviews the major methodological paradigms in VLA, including autoregressive-based modeling, diffusion-based approaches, reinforcement learning strategies, and hybrid or specialized designs. Fig. 3 presents a tree diagram illustrating the progression of these paradigms, with each branch highlighting representative works within the respective taxonomy. The taxonomy is organized chronologically, emphasizing how methodological innovations have progressively expanded the capabilities of VLA models.
3.1 Autoregression-Based Models in Vision Language Action Research
@@
Autoregression-based models constitute a classical yet effective paradigm for sequence generation in Vision–Language–Action (VLA) tasks. By treating action sequences as temporally dependent processes, these models generate actions step by step, conditioned on prior context, perceptual inputs, and task prompts. With the rapid advancement of Transformer architectures, recent VLA systems [29, 21, 30] have demonstrated the scalability and robustness of this approach. Representative works in these directions are summarized in Table I, collectively highlighting the versatility and generality of the autoregression-based modeling paradigm in VLA research.
@@ 3.1.3 Autoregressive Trajectory Generation and Visual Alignment Modeling
Autoregression-based methods have been adapted to diverse robot embodiments, from quadruped locomotion to bimanual manipulation, demonstrating the flexibility of vision–language–action frameworks [112, 69, 113]. Large-scale efforts such as OpenVLA [21, 114] further highlight cross-platform generalization and efficient adaptation, while latent motion token approaches [115, 71] point toward lightweight pretraining strategies.
@@ 3.1.5 Discussion
Innovations Autoregression-based models have driven major innovations in Vision–Language–Action research by unifying multimodal perception, language reasoning, and sequential action generation within scalable Transformer architectures. They support generalist agents capable of cross-task generalization, enable semantic planning through LLM integration, and extend trajectory generation to long-horizon and multimodal settings, structural optimizations such as token compression, parallel decoding, and quantization improve efficiency for real-world deployment.
@@
3.2 Diffusion-based Models in Vision Language Action Research
Diffusion (including flow matching, VAE, etc.) models have emerged as a transformative paradigm in generative artificial intelligence, showing remarkable potential within Vision Language Action (VLA) frameworks for embodied intelligence. In this subsection, we review the evolution of diffusion models in VLA systems, focusing on three critical dimensions. Representative works are summarized in Table II.
@@ 3.2 Diffusion-based Models in Vision Language Action Research
$D$ VLA [137] 2025 Modeled discretized action chunks with discrete diffusion and cross-entropy training.
@@
VQ-VLA [wang25vqvla] 2025 Employs vector-quantized tokenizers to reduce sim-to-real gaps.
@@ 3.2.3 Application Optimization and Deployment in Diffusion-Based VLA
The transition from laboratory prototypes to real-world deployment remains one of the most formidable challenges for diffusion-based Vision Language Action (VLA) systems. Meeting this challenge requires advances across three interconnected fronts: efficiency, adaptability, and robustness. Recent work demonstrates that rather than scaling model size indiscriminately, progress now hinges on principled optimization strategies, cognitive-inspired architectures, and practical deployment mechanisms.
Efficiency optimization has become a central theme. While diffusion models are resource-intensive, lightweight designs such as TinyVLA and SmolVLA [148, 149] show that pre-trained backbones with parameter-efficient tuning (e.g., LoRA) can cut training costs to single-GPU scales without sacrificing performance. Complementary strategies like VQ-VLA [wang25vqvla], which employs vector-quantized action tokenizers to narrow the sim-to-real gap, illustrate how efficiency gains can align with robustness. Collectively, these works reflect a paradigm shift toward “intelligent sparsity,” prioritizing performance-per-computation over brute-force scaling.
@@ 3.2.4 Discussion
3.3 Reinforcement-based Fine-Tune Models in Vision Language Action Research
@@ 3.3.1 Reinforcement-based Fine-Tune Strategies in VLA Research
Reinforcement-based Vision Language Action (VLA) methods integrate vision–language foundation models with reinforcement learning to enhance perception, reasoning, and decision-making. By leveraging visual and linguistic inputs, these methods generate context-aware actions within interactive and dynamic environments. They have emerged as a critical research direction for advancing autonomous driving, robotics, and broader embodied AI systems. Recent advances demonstrate that reinforcement-based VLA approaches can incorporate human feedback, adapt to novel tasks, and outperform purely supervised paradigms. The progression of these efforts is summarized in Table III.
Earlier methods improved robot manipulation skills using large-scale human video datasets or robot manipulation datasets by introducing reinforcement reward strategies [172, 173, 174, 181]. These methods aimed to investigate the promptability of pre-trained Vision Language Models (VLMs) in reinforcement learning, showing that even frozen models can support efficient downstream policy training through prompt embedding learning. VIP [172] derives a self-supervised goal-conditioned value function independent of actions, generating smooth embeddings that implicitly evaluate value via embedding distance.
@@ 3.4 Other Advanced Researches
TABLE IV: Evolution of Hybrid Architectures and Specialized Approacher in Vision Language Action Research.
@@ 3.4.3 Specialized Domain Adaptations and Applications
The VLA paradigm has also been extended to graphical user interface (GUI) interaction, where perception-action loops operate in fully digital spaces. ShowUI [245] adopts a vision–language–action pipeline for processing on-screen elements and generating control sequences for actions such as clicking, dragging, and form filling. Its strong performance on GUI-Bench underscores the applicability of VLA principles to non-physical manipulation tasks.
@@ 7 Challenges and Future Directions
7.1 Challenges of Vision Language Action Models
This section summarizes the open challenges and future directions in advancing Vision Language Action (VLA) models. Despite remarkable progress in recent years, the development of VLA models has gradually revealed critical bottlenecks. The most fundamental issue lies in the fact that current VLA systems are largely built upon the transfer of large-scale LLMs or VLMs. Although these models excel in semantic understanding and cross-modal alignment, they lack direct training and experience in interacting with the physical world. As a result, VLA systems often demonstrate the gap of “understanding the instruction but failing to execute the task” in real environments. This reflects a fundamental contradiction: the disconnection between semantic-level generalization and embodied capabilities in the physical world. How to achieve the transformation from non-embodied knowledge to embodied intelligence, and to truly bridge the gap between semantic reasoning and physical execution, remains the central challenge. Specifically, this contradiction is manifested in the following aspects.

### v3

# Pure Vision-Language-Action (VLA) Models: A Comprehensive Survey

@@ Abstract
The emergence of Vision-Language-Action (VLA) models marks a paradigm shift from traditional policy-based control to generalized robotics, reframing Vision-Language Models (VLMs) from passive sequence generators into active agents for manipulation and decision-making in complex, dynamic environments. This survey delves into advanced VLA methods, aiming to provide a clear taxonomy and a systematic, comprehensive review of existing research. It presents a comprehensive analysis of VLA applications across different scenarios and classifies VLA approaches into several paradigms: autoregression-based, diffusion-based, reinforcement-based, hybrid, and specialized methods; while examining their motivations, core strategies, and implementations in detail. In addition, foundational datasets, benchmarks, and simulation platforms are introduced. Building on the current VLA landscape, the review further proposes perspectives on key challenges and future directions to advance research in VLA models and generalizable robotics. By synthesizing insights from over three hundred recent studies, this survey maps the contours of this rapidly evolving field and highlights the opportunities and challenges that will shape the development of scalable, general-purpose VLA methods.
@@
Vision-Language-Action, Vision-Language Model, Robotics, Embodied AI
@@ 1 Introductions
To address these limitations, researchers have begun exploring the incorporation of large language models (LLMs) and vision-language models (VLMs) to enable more accurate and flexible robotic manipulation. Modern robotic manipulation methods [1, 2] typically leverage vision language generative paradigms (e.g., autoregressive models [3, 4, 5, 6] or diffusion models [7]), combined with large-scale datasets [8] and advanced fine-tuning strategies. We refer to these as VLA foundation models, which have substantially improved the quality of robotic manipulations. Fine-grained action control over generated content provides users with greater flexibility, unlocking the practical potential of VLA for task execution.
@@
Figure 3: Vision-Language-Action Taxonomy: From Autoregression-based, Diffusion-based, to Reinforcement-based and Hybrid/Specialized methods, multi-paradigm advances and practical adaptations in VLA. The taxonomy is organized along a chronological timeline.
@@ 2 Backgrounds
The emergence of Vision-Language-Action (VLA) models represents a significant advance toward general-purpose embodied intelligence. Traditional robotic systems typically rely on isolated perception pipelines, hand-engineered control strategies, or task-specific reinforcement learning. Although these approaches perform effectively in constrained environments, such as factory floors or laboratories, they generalize poorly to dynamic and unstructured settings. Modern robots can “see” using computer vision models, “understand” language through large language models, and “act” via controllers or learned policies; however, integrating these capabilities into a coherent and unified system remains a key challenge. VLA models address this challenge by providing a unified framework that grounds language in perception and maps it to executable actions.
@@ 2.1 Early: LLM/VLM fundamental Models
Against this backdrop, vision-language models (VLMs) emerged as a crucial bridge between unimodal learning and embodied intelligence. Early approaches (e.g., ViLBERT [16], VisualBERT [17]) aligned and fused images and text using dual-stream or single-stream Transformers, while contrastive learning methods (e.g., CLIP) mapped large-scale image–text pairs into a shared embedding space, enabling zero-shot and few-shot recognition and retrieval. More recently, instruction-tuned, dialogue-centric multimodal models (e.g., BLIP-2 [18], Flamingo [19], LLaVA [20]) have substantially strengthened open-ended cross-modal understanding, fine-grained grounding, and multi-turn reasoning, setting the stage for Vision-Language-Action (VLA) systems.
@@ 2.2.1 From LLM/VLM to VLA Models
Following this trajectory, research has naturally advanced toward VLA integration, which unifies visual perception, language understanding, and executable control within a single sequence modeling framework[21, 22, 1, 2, 23]. Typical designs encode images and instructions as prefix or context tokens, inject robot states and sensory feedback as state tokens, and autoregressively generate action tokens to produce control sequences, thereby closing the perception-language-action loop. Compared with traditional perception, planning, and control pipelines, VLA provides end-to-end cross-modal alignment and a unified treatment of goals, constraints, and intent. It inherits the semantic and instruction generalization of VLMs, while explicit state coupling and action generation confer robustness to environmental disturbances and long-horizon tasks. This progression—from unimodal to multimodal, and then to multimodal plus executable control—establishes the methodological foundation for systems that can not only see and understand, but also act.
@@ 2.2.2 The Supporting Role of Data and Simulation
The development of Vision-Language-Action (VLA) models in robotics relies heavily on high-quality datasets and realistic simulators capable of capturing the complexity of real-world scenarios. Recent robotic methods are typically deep learning–based and data-driven; consequently, dataset collection and annotation play a critical role in driving advancements in the field. Some datasets are collected in real-world settings, which requires substantial human effort and financial resources. To address these challenges, researchers also leverage large-scale human manipulation videos from the Internet as generalization datasets to provide auxiliary supervision for VLA model training. Despite these efforts, data collection remains costly, annotations are labor-intensive, and long-tail corner cases are often underrepresented. Other datasets are generated through robot simulators, which facilitate the collection of large-scale labeled data. Simulators provide diverse and controllable environments, flexible sensor configurations, realistic kinematic models, and interactive static and dynamic scenes, supporting both data collection and model evaluation. Representative datasets include Open X-Embodiment (OXE) [8], which integrates 22 robot datasets from 21 institutions, encompassing 527 skills and 160,266 tasks, and BridgeData [24], comprising 71 tasks across 10 environments in multiple domains. These resources standardize data formats, thereby promoting rapid development and reproducibility in VLA research. Simulators such as THOR [25], Habitat [26], MuJoCo [27], Isaac Gym [28], and CARLA [29] offer extensible virtual environments capable of generating multimodal annotations, including action trajectories, object states, and natural language instructions. Collectively, these datasets and simulation platforms mitigate the scarcity of real-world robot data and accelerate the training and evaluation of VLA models.
@@ 2.3 Future: Towards General Embodied Intelligence
VLA models occupy the forefront of research where vision, language, and action converge. They build upon breakthroughs in foundational models for perception and reasoning, emphasizing capabilities in human interaction and task execution, and extend these abilities to the physical world. By integrating the representational power of visual encoders, the reasoning capabilities of large language models, and the decision-making abilities of reinforcement learning and control frameworks, VLA models hold significant potential to bridge the “perception–understanding–action” gap. Despite facing challenges related to scalability, generalization, safety, and real-world deployment, VLA is widely recognized as a key frontier in embodied artificial intelligence. Although VLA has achieved notable success in Vision-Language-Action interactions, and benefits from advances in large-scale language models, it has not yet attained full generality in the domain of embodied intelligence. General embodied intelligence posits that human-like intelligent behavior relies not only on cognitive processing but also on the physical body, environmental perception, and feedback mechanisms, enabling interaction with the external world. To accommodate the demands of diverse tasks, general embodied intelligence can manifest through various types of robots, including humanoid robots for household applications, assembly robots with dexterous manipulators, and bionic robots with specialized capabilities. Clearly, general embodied intelligence has the potential to enable artificial systems to perform a wider array of tasks across diverse environments. VLA is currently evolving toward this vision of general embodied intelligence and holds considerable promise for realizing it.
3 Vision-Language-Action Models
@@
In recent years, Vision-Language-Action (VLA) models have undergone rapid and systematic development, driven by advances in multimodal representation learning, generative modeling, and reinforcement learning. To trace this evolution, this section reviews the major methodological paradigms in VLA, including autoregressive-based modeling, diffusion-based approaches, reinforcement learning strategies, and hybrid or specialized designs. Fig. 3 presents a tree diagram illustrating the progression of these paradigms, with each branch highlighting representative works within the respective taxonomy. The taxonomy is organized chronologically, emphasizing how methodological innovations have progressively expanded the capabilities of VLA models.
3.1 Autoregression-Based Models in Vision-Language-Action Research
@@
Autoregression-based models constitute a classical yet effective paradigm for sequence generation in Vision-Language-Action (VLA) tasks. By treating action sequences as temporally dependent processes, these models generate actions step by step, conditioned on prior context, perceptual inputs, and task prompts. With the rapid advancement of Transformer architectures, recent VLA systems [30, 21, 31] have demonstrated the scalability and robustness of this approach. Representative works in these directions are summarized in Table I, collectively highlighting the versatility and generality of the autoregression-based modeling paradigm in VLA research.
@@ 3.1.3 Autoregressive Trajectory Generation and Visual Alignment Modeling
Autoregression-based methods have been adapted to diverse robot embodiments, from quadruped locomotion to bimanual manipulation, demonstrating the flexibility of Vision-Language-Action frameworks [113, 70, 114]. Large-scale efforts such as OpenVLA [21, 115] further highlight cross-platform generalization and efficient adaptation, while latent motion token approaches [116, 72] point toward lightweight pretraining strategies.
@@ 3.1.5 Discussion
Innovations Autoregression-based models have driven major innovations in Vision-Language-Action research by unifying multimodal perception, language reasoning, and sequential action generation within scalable Transformer architectures. They support generalist agents capable of cross-task generalization, enable semantic planning through LLM integration, and extend trajectory generation to long-horizon and multimodal settings, structural optimizations such as token compression, parallel decoding, and quantization improve efficiency for real-world deployment.
@@
3.2 Diffusion-based Models in Vision-Language-Action Research
Diffusion (including flow matching, VAE, etc.) models have emerged as a transformative paradigm in generative artificial intelligence, showing remarkable potential within Vision-Language-Action (VLA) frameworks for embodied intelligence. In this subsection, we review the evolution of diffusion models in VLA systems, focusing on three critical dimensions. Representative works are summarized in Table II.
@@ 3.2 Diffusion-based Models in Vision-Language-Action Research
DD VLA [138] 2025 Modeled discretized action chunks with discrete diffusion and cross-entropy training.
@@
VQ-VLA [151] 2025 Employs vector-quantized tokenizers to reduce sim-to-real gaps.
@@ 3.2.3 Application Optimization and Deployment in Diffusion-Based VLA
The transition from laboratory prototypes to real-world deployment remains one of the most formidable challenges for diffusion-based Vision-Language-Action (VLA) systems. Meeting this challenge requires advances across three interconnected fronts: efficiency, adaptability, and robustness. Recent work demonstrates that rather than scaling model size indiscriminately, progress now hinges on principled optimization strategies, cognitive-inspired architectures, and practical deployment mechanisms.
Efficiency optimization has become a central theme. While diffusion models are resource-intensive, lightweight designs such as TinyVLA and SmolVLA [149, 150] show that pre-trained backbones with parameter-efficient tuning (e.g., LoRA) can cut training costs to single-GPU scales without sacrificing performance. Complementary strategies like VQ-VLA [151], which employs vector-quantized action tokenizers to narrow the sim-to-real gap, illustrate how efficiency gains can align with robustness. Collectively, these works reflect a paradigm shift toward “intelligent sparsity,” prioritizing performance-per-computation over brute-force scaling.
@@ 3.2.4 Discussion
3.3 Reinforcement-based Fine-Tune Models in Vision-Language-Action Research
@@ 3.3.1 Reinforcement-based Fine-Tune Strategies in VLA Research
Reinforcement-based Vision-Language-Action (VLA) methods integrate vision–language foundation models with reinforcement learning to enhance perception, reasoning, and decision-making. By leveraging visual and linguistic inputs, these methods generate context-aware actions within interactive and dynamic environments. They have emerged as a critical research direction for advancing autonomous driving, robotics, and broader embodied AI systems. Recent advances demonstrate that reinforcement-based VLA approaches can incorporate human feedback, adapt to novel tasks, and outperform purely supervised paradigms. The progression of these efforts is summarized in Table III.
Earlier methods improved robot manipulation skills using large-scale human video datasets or robot manipulation datasets by introducing reinforcement reward strategies [174, 175, 176, 183]. These methods aimed to investigate the promptability of pre-trained Vision-Language Models (VLMs) in reinforcement learning, showing that even frozen models can support efficient downstream policy training through prompt embedding learning. VIP [174] derives a self-supervised goal-conditioned value function independent of actions, generating smooth embeddings that implicitly evaluate value via embedding distance.
@@ 3.4 Other Advanced Researches
TABLE IV: Evolution of Hybrid Architectures and Specialized Approacher in Vision-Language-Action Research.
@@ 3.4.3 Specialized Domain Adaptations and Applications
The VLA paradigm has also been extended to graphical user interface (GUI) interaction, where perception-action loops operate in fully digital spaces. ShowUI [247] adopts a Vision-Language-Action pipeline for processing on-screen elements and generating control sequences for actions such as clicking, dragging, and form filling. Its strong performance on GUI-Bench underscores the applicability of VLA principles to non-physical manipulation tasks.
@@ 7 Challenges and Future Directions
7.1 Challenges of Vision-Language-Action Models
This section summarizes the open challenges and future directions in advancing Vision-Language-Action (VLA) models. Despite remarkable progress in recent years, the development of VLA models has gradually revealed critical bottlenecks. The most fundamental issue lies in the fact that current VLA systems are largely built upon the transfer of large-scale LLMs or VLMs. Although these models excel in semantic understanding and cross-modal alignment, they lack direct training and experience in interacting with the physical world. As a result, VLA systems often demonstrate the gap of “understanding the instruction but failing to execute the task” in real environments. This reflects a fundamental contradiction: the disconnection between semantic-level generalization and embodied capabilities in the physical world. How to achieve the transformation from non-embodied knowledge to embodied intelligence, and to truly bridge the gap between semantic reasoning and physical execution, remains the central challenge. Specifically, this contradiction is manifested in the following aspects.